Boxing gloves are padded to lessen the force of a blow. (a) Calculate the force exerted by a boxing glove on an opponent’s face, if the glove and face compress 7.50 cm during a blow in which the 7.00-kg arm and glove are brought to rest from an initial speed of 10.0 m/s. (b) Calculate the force exerted by an identical blow in the gory old days when no gloves were used and the knuckles and face would compress only 2.00 cm. (c) Discuss the magnitude of the force with glove on. Does it seem high enough to cause damage even though it is lower than the force with no glove?
Answer #1

A)

Force, F = mu^2/2x

F = (7 x 10^2)/(2 x 0.075) = 4666.67 N

B)

Force, F2 = (7 x 10^2)/(2 x 0.02) = 17500 N

C)

Force is equal to 477 kg mass. It is enough to cause damage on the face even it is lower than force with no glove

Comment in case any doubt please rate my answer.....